OpenStudy (anonymous):

Find the period of \[f(x)) = 4 sin^2x + 7 cos^2 x\]

OpenStudy (anonymous):

My guess is \(\pi\).

OpenStudy (anonymous):

Yup...@Ishaan94 u r correct

OpenStudy (anonymous):

Can u ....Solve this....@Ishaan94

OpenStudy (anonymous):

Period of \(b\sin^2x + a\cos^2x\) is same as \(c\sin^2x\). Use \(\sin^2\theta + \cos^2\theta = 1\).

OpenStudy (anonymous):

Did nt get u @Ishaan94

OpenStudy (anonymous):

\[4\sin^2x + 7\cos^2x = 4 - 11 \cos^2x\] Since \(\cos ^2x\)'s period \(\pi\), \(4 - 11 \cos^2x\) would \(\pi\) as well.

OpenStudy (anonymous):

... would be* ...

OpenStudy (anonymous):

Yes...thxxxx